# Degree two Gopakumar–Vafa invariants of local curves

### v3

@@ 3.1.7. $\lambda=(3,1^{3})$
Consider $Z$, the subvariety of pairs $(x,(y,z,w))\in C\times\operatorname{Sym}^{3}(C)$ giving solutions to $\mathcal{O}_{C}(3x+y+z+w)\cong L^{\otimes 2}$. Projecting along the morphism that forgets $y,z,w$, this is a $\mathbb{P}^{1}$-bundle $\pi\colon Z\rightarrow C$, and so has Euler characteristic $-4=2*(-2)$.
@@ 3.1.8. $\lambda=(2^{3})$
[…]
@@ 3.1.9. $\lambda=(2,2,1,1)$
Consider the rational map $\phi\colon C\times C\dashrightarrow\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})$ which sends generic $(x,y)\in C\times C$ to the one-dimensional subspace $\operatorname{H}^{0}(L^{\otimes 2}(-2x-2y))\subset\operatorname{H}^{0}(L^{\otimes 2})$. As discussed above, the indeterminancy locus of this rational map is $32$ points in $C\times C$. In the following, we calculate the line bundle inducing the rational map $\phi$. Consider the following diagram:
@@
where $p_{12}\colon C\times(C^{\times 2})\to C\times C$ is given by forgetting the third factor, and $\Delta\subset C\times C$ is the diagonal. The other parts of the calculations for $*_{2}$ are similar.
@@ 3.1.10. $\lambda=(2,1^{4})$
$e(P)=3*(-2)=-6.$
Let $D$ be the pull back of $H_{\infty}$ to $Z$, and $F$ a fibre of $r$. As in the case $\lambda=(3,1,1,1)$, we can see that $D\cap F$ is isomorphic to $\mathbb{P}^{1}$, linearly embedded into $F$. In other words, $D$ is a $\mathbb{P}^{1}$-bundle over $C$ and so $e(D)=-4=(-2)*2$.
@@ 3.2. A degeneracy locus calculation
We have $e(Z)=c_{2}(\mathcal{F})c_{1}(C^{\times 3})-c_{1}(\mathcal{F})c_{2}(\mathcal{F})-c_{3}(\mathcal{F})$.
@@
$e(Z)$ $=s_{2}(V\otimes\mathcal{O}-\mathcal{F}^{\vee})c_{1}(C^{\times 3})-s_{(2,1)}(V\otimes\mathcal{O}-\mathcal{F}^{\vee})-2s_{3}(V\otimes\mathcal{O}-\mathcal{F}^{\vee})$
@@
We denote by $*_{i}$ the $i$th component of $\operatorname{ch}(\sigma^{*}L^{\otimes 2}\otimes\mathcal{O}_{3\Delta_{1}+2\Delta_{2}+\Delta_{3}}).\operatorname{td}_{\mu}$. It is obvious that
$*_{0}=0,\quad*_{1}=3\Delta_{1}+2\Delta_{2}+\Delta_{3}.$
@@
By pushing forward $*_{i}$ via $\mu$, we obtain the result. ∎
@@
$e(Z)$ $=c_{2}(\mathcal{F})c_{1}(C^{\times 3})-c_{1}(\mathcal{F})c_{2}(\mathcal{F})-c_{3}(\mathcal{F})$
@@ 3.3. Fibrewise GV invariants
$g=3$ $4$ $5$ $6$ $\chi(U_{\epsilon}\cap S_{\lambda})$
$\lambda=(5,1)$ 00 $3$ $-4$ $1$ $50$
$(4,2)$ $-1$ $4$ $-4$ $1$ $128$
$(4,1,1)$ 00 $1$ $-3$ $1$ $-216$
$(3,3)$ 00 $4$ $-4$ $1$ $81$
$(3,2,1)$ 00 $2$ $-3$ $1$ $-668$
$(3,1^{3})$ 00 00 $-2$ $1$ $542$
$(2^{3})$ $-1$ $3$ $-3$ $1$ $-128$
$(2,2,1,1)$ 00 $1$ $-2$ $1$ $968$
$(2,1^{4})$ 00 00 $-1$ $1$ $-1012$
$(1^{6})$ 00 00 00 $1$ $256$
@@
where we denote by $r(\lambda)$ the number of rows of $\lambda$. We give a little more detail, firstly in order to explain the sign $(-1)^{g}$ appearing in the LHS of (3.4). Note that $102g(C_{a})-2=$ 10 for all $a$. By [MS13, MY14], we have the following (cf. [MT18][Equation (4.4)]):
$\sum_{l}\chi(\operatorname{IC}_{C_{a}^{[l]}})q^{l}=q^{g(C_{a})-1}\cdot\frac{1}{(q^{-1/2}+q^{1/2})^{2}}\sum_{g}\chi(\operatorname{Gr}_{P}^{i}\operatorname{IC}(\overline{J}))q^{i}.$
@@ 3.4. Contributions from the nearby hyperplane
$n_{g}(U_{\varepsilon})=\sum_{\lambda\vdash 6}n_{g}(\lambda)\cdot e(U_{\varepsilon}\cap S_{\lambda}).$
@@
$n_{3}(U_{\varepsilon})$ $=-1\cdot e(U_{\varepsilon}\cap S_{(4,2)})-1\cdot e(U_{\varepsilon}\cap S_{(2,2,2)})$
@@ 3.5. GV invariants for twisted Higgs bundles
Let $\hat{M}$ denote the moduli space of $L$-twisted $\operatorname{PGL}_{n}$-Higgs bundles.
We first describe a basis of $\operatorname{H}^{i}(\hat{M})$ for each $i$. By [Moz12, Subsection 6.3], adopting the notation of loc. cit., the Poincare polynomial of $\hat{M}$ is
@@
On the other hand, the cohomology ring $\operatorname{H}^{*}(\hat{M})$ has the following multiplicative generators [dCHM12, HT03, HT04]:
@@
Using the relations in [dCHM12, Theorem 1.2.10, Proposition 4.2.1] and the equation (3.5), we have that
@@
where we put $\gamma\coloneqq-2(\psi_{1}\psi_{3}+\psi_{2}\psi_{4})$. In the following, we only explain the above equalities for $\operatorname{H}^{7}(\hat{M})$ and $\operatorname{H}^{8}(\hat{M})$. The arguments are similar and simpler for lower cohomological degrees. For $\operatorname{H}^{7}(\hat{M})$, we have
@@
where the second equality follows from [dCHM12, Proposition 4.2.1]. On the other hand, by (3.5), we also know that $\dim\operatorname{H}^{7}(\hat{M})=4$. Hence, we can conclude that $\{\alpha^{2}\psi_{i}\}_{1\leq i\leq 4}$ is a basis of $\operatorname{H}^{7}(\hat{M})$. For $\operatorname{H}^{8}(\hat{M})$, we have
(3.6) $\operatorname{H}^{8}(\hat{M})$ $=\langle\alpha^{4},\alpha^{2}\beta,\alpha\psi_{i}\psi_{j},\beta^{2}\rangle_{1\leq i<j\leq 4}$
@@
where $I^{a}_{b}\subset\mathbb{Q}[\alpha,\beta,\gamma]$ is an ideal defined in [dCHM12, Definition 1.2.8] for $a,b\geq 0$. Indeed, we have
@@
with $\Lambda^{2}_{0}$ as defined in loc. cit., we conclude that the elements in (3.7), multiplied by $\alpha\in I^{0}_{3}$, are all zero in $\operatorname{H}^{*}(\hat{M})$. By the above arguments, the third equality in (3.6) holds.
Now, we rearrange the above basis of $\operatorname{H}^{*}(\hat{M})$ using the perverse degree instead of the cohomological degree. Note that the generators $\alpha,\beta,\psi_{i}$ all have perverse degree $2$ by [dCHM12, Theorem 4.2.2]. Hence, we obtain
@@
Here we have followed the notation and normalisation conventions of [dCHM12], regarding the perverse filtration on $\operatorname{H}^{*}(\hat{M})$. Precisely:
@@ 4. Higher genus
In this section, we fix a smooth projective curve $C$ of genus $g(C)\geq 3$ and a line bundle $L$ of degree $12g(C)-$ 1.
@@
where $U\subset\widetilde{B}_{L}$ is the dense open subset parametrising smooth spectral curves and $\pi\colon\mathcal{C}\to U$ is the universal curve. Since a smooth spectral curve has genus $24g(C)-$ 2, the degree of the Laurent polynomial $\sum_{i\in\mathbb{Z}}\chi({}^{p}\mathcal{H}^{i}(\mathbf{R}h_{*}\operatorname{IC}_{M}))q^{i}$ is less than or equal to $24g(C)-$ 2. The coefficient of $q^{4g(C)-2}$ is
@@
$\operatorname{H}^{*}(M)\cong\operatorname{H}^{*}(\operatorname{Pic}^{0}(C))\otimes\operatorname{H}^{*}(\hat{M}).$
@@
From this, we can see that $n_{g,2[C]}(\operatorname{Tot}_{C}(L))=0$ for $g\leq g(C)-1$. Furthermore, we have $n_{g,2[C]}(\operatorname{Tot}_{C}(L))=-\chi(\hat{M})$. This can be computed as follows: Let $P_{t}(M)$ denote the Poincare polynomial of $M$. We can compute $P_{-t}(M)$ by substituting $u=v=-t$ in [CDP11, Equation (A.7)]:
@@
    $n_{2g(C)-1}(U_{\varepsilon})=-\sum_{\lambda}e(S_{\lambda}\cap U_{\varepsilon})$, where $\lambda=(\lambda_{i})$ runs over all partitions of $24g(C)-$ 2 such that $\lambda_{i}$ is even for every $i$.
@@
[…]
$-2\chi(C)-\deg(L^{\otimes 2})=8g(C)-6$
@@
[…]
(4.1) $-2\chi(C)+\#\{i\colon\lambda_{i}\text{ is odd}\}$
@@
Let $L\in\operatorname{Pic}^{2g(C)-1}(C)$ and $[N]\in\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})$ be arbitrary. Then we have

### v4

@@ 3.1.7. $\lambda=(3,1^{3})$
Consider $Z$, the subvariety of pairs $(x,(y,z,w))\in C\times\operatorname{Sym}^{3}(C)$ giving solutions to $\mathcal{O}_{C}(3x+y+z+w)\cong L^{\otimes 2}$. Projecting along the morphism that forgets $y,z,w$, this is a $\mathbb{P}^{1}$-bundle $\pi\colon Z\rightarrow C$, and so has Euler characteristic 4=2(2)-4=2*(-2).
@@ 3.1.8. $\lambda=(2^{3})$
$\operatorname{Sym}^{3}(C)$$\operatorname{Sym}^{6}(C)$$\operatorname{Pic}^{3}(C)$$\operatorname{Pic}^{6}(C).$$\scriptstyle{2(-)}$$\scriptstyle{\operatorname{AJ}^{3}}$$\scriptstyle{\operatorname{AJ}^{6}}$$\scriptstyle{(-)^{\otimes 2}}$
@@ 3.1.9. $\lambda=(2,2,1,1)$
Consider the rational map $\phi\colon C\times C\dashrightarrow\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})$ which sends generic $(x,y)\in C\times C$ to the one-dimensional subspace $\operatorname{H}^{0}(L^{\otimes 2}(-2x-2y))\subset\operatorname{H}^{0}(L^{\otimes 2})$. As discussed above, the indeterminacy locus of this rational map is $32$ points in $C\times C$. In the following, we calculate the line bundle inducing the rational map $\phi$. Consider the diagram:
@@
where $p_{12}\colon C\times(C^{\times 2})\to C\times C$ is given by forgetting the third factor, and $\Delta\subset C\times C$ is the diagonal. The other parts of the calculations for $_{2}$ are similar.
@@ 3.1.10. $\lambda=(2,1^{4})$
$\chi(P)=3*(-2)=-6.$
Let $D$ be the pull back of $H_{\infty}$ to $Z$, and $F$ a fibre of $r$. As in the case $\lambda=(3,1,1,1)$, we can see that $D\cap F$ is isomorphic to $\mathbb{P}^{1}$, linearly embedded into $F$. In other words, $D$ is a $\mathbb{P}^{1}$-bundle over $C$ and so $\chi(D)=-4=(-2)*2$.
@@ 3.2. A degeneracy locus calculation
We have $\chi(Z)=c_{2}(\mathcal{F})c_{1}(C^{\times 3})-c_{1}(\mathcal{F})c_{2}(\mathcal{F})-c_{3}(\mathcal{F})$.
@@
$\chi(Z)$ $=s_{2}(V\otimes\mathcal{O}-\mathcal{F}^{\vee})c_{1}(C^{\times 3})-s_{(2,1)}(V\otimes\mathcal{O}-\mathcal{F}^{\vee})-2s_{3}(V\otimes\mathcal{O}-\mathcal{F}^{\vee})$
@@
We denote by $_{i}$ the $i$th component of $\operatorname{ch}(\sigma^{*}L^{\otimes 2}\otimes\mathcal{O}_{3\Delta_{1}+2\Delta_{2}+\Delta_{3}}).\operatorname{td}_{\mu}$. It is obvious that
$_{0}=0,\quad*_{1}=3\Delta_{1}+2\Delta_{2}+\Delta_{3}.$
@@
By pushing forward $_{i}$ via $\mu$, we obtain the result. ∎
@@
$\chi(Z)$ $=c_{2}(\mathcal{F})c_{1}(C^{\times 3})-c_{1}(\mathcal{F})c_{2}(\mathcal{F})-c_{3}(\mathcal{F})$
@@ 3.3. Fibrewise GV invariants
$\lambda$ $g=3$ 44 55 66 $\chi(U_{\epsilon}\cap S_{\lambda})$
$(5,1)$ 0 33 4-4 11 $50$
$(4,2)$ 1-1 44 4-4 11 $128$
$(4,1,1)$ 0 11 3-3 11 216-216
$(3,3)$ 0 44 4-4 11 $81$
$(3,2,1)$ 0 22 3-3 11 668-668
$(3,1^{3})$ 0 0 2-2 11 $542$
$(2^{3})$ 1-1 33 3-3 11 128-128
$(2,2,1,1)$ 0 11 2-2 11 $968$
$(2,1^{4})$ 0 0 1-1 11 1012-1012
$(1^{6})$ 0 0 0 11 $256$
@@
where we denote by $r(\lambda)$ the number of rows of $\lambda$. We give a little more detail, firstly in order to explain the sign $(-1)^{g}$ appearing in the LHS of (3.4). Note that $2g(C_{a})-2=10$ for all $a$. By [MS13, MY14], we have the following (cf. [MT18][Equation (4.4)]):
$\sum_{l}\chi(\operatorname{IC}_{C_{a}^{[l]}})q^{l}=q^{g(C_{a})-1}\cdot\frac{1}{(q^{-1/2}+q^{1/2})^{2}}\sum_{i}\chi(\operatorname{Gr}_{P}^{i}\operatorname{IC}(\overline{J}))q^{i}.$
@@ 3.4. Contributions from the nearby hyperplane
$n_{g}(U_{\varepsilon})=\sum_{\lambda\vdash 6}n_{g}(\lambda)\cdot\chi(U_{\varepsilon}\cap S_{\lambda}).$
@@
$n_{3}(U_{\varepsilon})$ $=-1\cdot\chi(U_{\varepsilon}\cap S_{(4,2)})-1\cdot\chi(U_{\varepsilon}\cap S_{(2,2,2)})$
@@ 3.5. GV invariants for twisted Higgs bundles
Let $\hat{M}$ denote the moduli space of semistable $L$-twisted $\operatorname{PGL}_{n}$-Higgs bundles.
We first describe a basis $\operatorname{H}^{i}(\hat{M}) fo$r each $i$. By [MOZ12, Subsection 6.3], adopting the notation of loc. cit., the Poincare polynomial of $\hat{M}$ is
@@
[…]
@@
Using the relations in [dHM12, Theorem 1.2.10, Proposition 4.2.1] and the equation (3.5), we have that
@@
[…]
@@
[…]
@@
where $I^{a}_{b}\subset\mathbb{Q}[\alpha,\beta,\gamma]$ is an ideal defined in [dHM12, Definition 1.2.8] for $a,b\geq 0$. Indeed, we have
@@
[…]
@@
[…]
@@ 4. Higher genus
In this section, we fix a smooth projective curve $C$ of genus $g(C)\geq 3$ and a line bundle $L$ of degree 2g(C)12g(C)-1.
@@
where $U\subset\widetilde{B}_{L}$ is the dense open subset parametrising smooth spectral curves and $\pi\colon\mathcal{C}\to U$ is the universal curve. Since a smooth spectral curve has genus 4g(C)24g(C)-2, the degree of the Laurent polynomial $\sum_{i\in\mathbb{Z}}\chi({}^{p}\mathcal{H}^{i}(\mathbf{R}h_{*}\operatorname{IC}_{M}))q^{i}$ is less than or equal to 4g(C)24g(C)-2. The coefficient of $q^{4g(C)-2}$ is
@@
[…]
@@
[…]
@@
    $n_{2g(C)-1}(U_{\varepsilon})=-\sum_{\lambda}\chi(S_{\lambda}\cap U_{\varepsilon})$, where $\lambda=(\lambda_{i})$ runs over all partitions of 4g(C)24g(C)-2 such that $\lambda_{i}$ is even for every $i$.
@@
Given a partition $\lambda$ of 4g(C)24g(C)-2 and an element $a\in S_{\lambda}\cap U_{\varepsilon}$, let $C_{a}\to C$ be the corresponding spectral curve. In view of the recursion formula (3.3) computing $P(T_{2,i})$ and the formula (3.4) for $\chi(\operatorname{Hilb}^{n}(C_{a}))$, we see that the polynomial $\int_{C_{a}^{[*]}}q^{2l}d\chi$ has degree
$2\chi(C)-\deg(L^{\otimes 2})=8g(C)-6$
@@
On the other hand, if we write $\int_{C_{a}^{[*]}}q^{2l}d\chi$ as a polynomial in $(q^{-1}-q)$, the lowest term has degree
(4.1) $2\chi(C)+\#\{i\colon\lambda_{i}\text{ is odd}\}$
@@
Let $L\in\operatorname{Pic}^{2g(C)-1}(C)$ and $N]\in\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})$ be arbitrary. Then we have